OpenStudy (the_puppet1):

Event A: Rolling a number less than 4 on a number cube Event B: Flipping H on a coin toss P(A) = ___/____ P(B) = ___/____ P(A and B) = ____/____

OpenStudy (peachpi):

P(A) = # of numbers less than 4 / # of sides of a cube P(B) = # of H on a coin / # of sides on a coin P(A and B) = P(A) * P(B)

OpenStudy (the_puppet1):

Plz this is so confusing! For meh! @peachpi I had a ruff week

OpenStudy (peachpi):

For P(A) answer two questions. 1. How many numbers on a cube are less than 4? 2. How many sides are there on a cube?

OpenStudy (the_puppet1):

3...

OpenStudy (peachpi):

how many sides are there on a cube?

OpenStudy (the_puppet1):

6 on a cube

OpenStudy (peachpi):

So P(A) = 3/6 = 1/2

OpenStudy (peachpi):

For P(B) 1. How many heads on a coin? 2. How many sides on a coin?

OpenStudy (the_puppet1):

1/0?

OpenStudy (peachpi):

there are two sides on a coin, heads and tails

OpenStudy (peachpi):

so P(B) = ?

OpenStudy (the_puppet1):

P(B) = 1/2

OpenStudy (peachpi):

yes

OpenStudy (peachpi):

A and B are independent events, one doesn't affect the other, so P(A and B) = P(A) * P(B)

OpenStudy (the_puppet1):

So P(A and B) = 1/2 * 1/2?

OpenStudy (peachpi):

right, and what is that?

OpenStudy (the_puppet1):

= 0.25

OpenStudy (peachpi):

yes

OpenStudy (the_puppet1):

So it's 2/5?

OpenStudy (peachpi):

no it's 0.25

OpenStudy (peachpi):

0.25 in fraction is 1/4

OpenStudy (the_puppet1):

Thank you

OpenStudy (peachpi):

you're welcome